3[3-3(x+1)]+1=2(-14-x)+2x+14
Answer by atif.muhammad(135)

3[3-3(x+1)]+1=2(-14-x)+2x+14
3[3 - 3x - 3] + 1 = - 28 - 2x + 2x + 14
3[3x] + 1 = - 14 --> forgot the -ve sign in front of the 3x term (LongJonSilver)
9x + 1 = -14
9x = -15
3x = -5
x = -5/3